# Essential Dimension of Central Simple Algebras when the Characteristic is Bad

### v1

@@ 9. Algebras of prime degree
Fix a prime degree $p$ and a field $F$ of $\operatorname{char}(F)=p$. A cyclic algebra of degree $p$ over $F$ is generated by two elements $i$ and $j$ such that $i^{p}-i=\alpha$ for some $\alpha\in F$, $j^{p}=\beta$ for some $\beta\in F^{\times}$ and $jij^{-1}=i+1$. This algebra is denoted by the symbol (also called a Hilbert symbol) $\alpha,\beta)_{p,F}$. The extension $F[i]/F$ is either a cyclic field extension of degree $p$ with Galois group $\langle\sigma\rangle$ where $\sigma(i)=i+1$ or $F\times\dots\times F$, an étale extension. In both cases, we can define the norm $N_{\alpha}:F[i]\rightarrow F$ by $x\mapsto x\sigma(x)\dots\sigma^{p-1}(x)$ where $\sigma(i)=i+1$. The algebra $\alpha,\beta)_{p,F}$ is “split”, i.e., isomorphic to $M_{p}(F)$, when $\beta=N_{\alpha}(y)$ for some $y\in F[i]$. If no such $y$ exists, then $\alpha,\beta)_{p,F}$ is a division algebra. Different symbols can represent the same isomorphism class. In particular $\alpha,\beta)_{p,F}=[\alpha+\lambda^{p}-\lambda+x^{p}\beta^{n},\beta)_{p,F}$ for any choice of $\lambda,x\in F$ and $n\in\{1,\dots,p-1\}$, and $\alpha,\beta)_{p,F}=[\alpha,\beta N_{\alpha}(y))_{p,F}$ for any $y\in F[i]$ (note that we use == also for $\cong$ as is common in the literature).
It is still unknown whether every algebra of degree $p$ over $F$ is cyclic. This is known when $p=2$ or $3$, but open for $p\geqslant 5$. In the case of $p$-special fields, however, this is known to be true. Recall that a field $F$ is $p$-special, if it has no finite extension fields of degree prime to $p$ [EKM08, Section 101.B]. Note that by citeLotscher:2013, all finite extensions of a $p$-special field are of degree $p^{m}$ for some positive integer $m$.
Theorem 9.1 ([Alb68]).
@@
This is an immediate result of Theorem [Alb68, Chapter IV, Theorem 32]. However, since there is an argument in more modern terms, we prefer to include it here.
Consider $A$ of degree $p$ over $F$. A necessary and sufficient condition for $A$ to be cyclic is the inclusion of a non-central element $y$ satisfying $y^{p}\in F$ ([Alb68, Lemma 10]). Recall that there are the characteristic coefficients $f_{1},\dots,f_{p-1},f_{p}$ from $A$ to $F$, and that each $x\in A$ satisfies $x^{p}-f_{1}(x)x^{p-1}-\dots-f_{p-1}(x)x-f_{p}(x)=0$. It is therefore enough to find a nontrivial solution to the system $f_{1}(x)=f_{2}(x)=\dots=f_{p-1}(x)=0$ over a subspace of $A$ consisting of noncentral elements, and then the solution $y$ will be a noncentral element satisfying $y^{p}\in F$. Since the center of $A$ is a subspace of dimension 1, there exists a subspace $V$ (satisfying $A=V\oplus F$) consisting of noncentral elements (apart from zero) of dimension $p^{2}-1$. By the Nullstellensatz for $p$-special fields ([Pfi95, Theorem 1.13]), there exists a nontrivial solution to the system $f_{1}(x)=f_{2}(x)=\dots=f_{p-1}(x)=0$ over $V$, and that settles the issue. ∎
Corollary 9.2 ([RY00, Lemma 8.5 (7)]).
The essential $p$-dimension of $\operatorname{Alg}_{p,p}$ is $2$ when the algebraically closed field $k$ over which this functor is defined is of characteristic $p$.
@@
It cannot be greater than 2 because of the previous theorem. In order to show that it is exactly 2, it is enough to point out one cyclic division algebra of degree $p$, because the transcendence degree bounds the cohomological $p$-dimension, and for the existence of division cyclic algebras, it must be at least 2. By [CC17, Remark] (using valuation theory, see [TW15] for background), the generic cyclic algebra $\alpha,\beta)_{p,F}$ is division for $F=k(\alpha,\beta)$. Since it remains division under prime-to-$p$ extension, the statement readily follows.
Remark 9.3.
It is important to note that in [Pfi95], a $p$-special field is defined as a field which has only field extensions of degrees powers of $p$. The definition given a few paragraphs above sounds more inclusive than the first, but in reality they coincide (see [LMMR13, Lemma 2.1]). In [LMMR13] they also explain how a $p$-special field is obtained as the closure under extensions to all prime-to-$p$ extensions, and thus, the relation between the $p$-special closure and the essential $p$-dimension is obvious.
@@ 10. Algebras of Prime Exponent
By [GS06, Theorem 9.1.4], which dates back to Teichmüller, ${{}_{p}Br}(F)$ is generated by cyclic algebras of degree $p$ for fields $F$ of $\operatorname{char}(F)=p>0$. Tensor products of cyclic algebras of degree $p$ therefore have a special place in the theory of central simple algebras.
@@
As mentioned in Section 8, there is a description of ${{}_{p}Br}(F)$ in terms of differential forms. Recall that $\Omega_{F}$ is the ring of differential forms over $F$, i.e., sums of terms of the form adbadb, where $a,b\in F$, and $d(bc)=cdb+bdc$. The group ${{}_{p}Br}(F)$ is isomorphic to the cokernel of $\wp$, where $\wp:\Omega_{F}\rightarrow\Omega_{F}/dF$ is given by $a\frac{db}{b}\mapsto(a^{p}-a)\frac{db}{b}$. The isomorphism takes $a\frac{db}{b}$ to $a,b)_{p,F}$ and, as the argument in Section 8 with $n=1$ shows, the $p$-rank of $F$ bounds the symbol length of ${{}_{p}Br}(F)$, i.e., the maximal number of cyclic algebras of degree $p$ required in order to express any given class in ${{}_{p}Br}(F)$ is bounded from above by the $p$-rank of $F$.
@@
The fact that $\operatorname{ed}((\mathbb{Z}/p\mathbb{Z})^{\times n})$ is 1 over fields of characteristic $p$ ([Led04]) implies that the essential dimension of tensor products of $n$ cyclic algebras of degree $p$ is bounded from above by $n+1$. Another way to see it, without using [Led04], is the observation made in [CQM24] that every tensor product of $n$ cyclic algebras of degree $p$ can be written as
(1) $[\alpha_{1},\alpha_{2})_{p,F}\otimes[\alpha_{2},\alpha_{3})_{p,F}\otimes\dots\otimes[\alpha_{n},\alpha_{n+1})_{p,F}.$
@@
This observation is based on the rule $\alpha,\beta)_{p,F}=[\alpha(\beta+x^{p})\beta^{-1},\beta+x^{p})_{p,F}$, which means that each $\alpha,\beta)_{p,F}\otimes[\gamma,\delta)_{p,F}$ can be modified into $\alpha(\beta+x^{p})\beta^{-1},\beta+x^{p})_{p,F}\otimes[\gamma+x^{p}-x,\delta)_{p,F}$ where $x=\gamma-\beta$ (i.e., $\gamma+x^{p}-x=\beta+x^{p}$).
On the other hand, the fact that the generic tensor product $\alpha_{1},\beta_{1})_{p,F}\otimes\dots\otimes[\alpha_{n},\beta_{n})_{F,p}$ over $F=k(\alpha_{1},\beta_{1},\dots,\alpha_{n},\beta_{n})$ is of essential dimension $n+1$ ([McK17]) and it can also be written in the form (1), means that the essential dimension of $\alpha_{1},\alpha_{2})_{p,F}\otimes[\alpha_{2},\alpha_{3})_{p,F}\otimes\dots\otimes[\alpha_{n},\alpha_{n+1})_{p,F}$ over $F=k(\alpha_{1},\dots,\alpha_{n},\alpha_{n+1})$ is also $n+1$. Note that the latter is a division algebra (see [Cha20]).
@@
All this means that $\operatorname{ed}(\operatorname{Alg}_{p^{n},p})\geqslant n+1$. For $p=2$ and $n=3$, this improves the lower bound produced in [Bae11] from 3 to 4. There is another argument for this lower bound, that appeared in [CM20]. The point is that unless $p=n=2$, indecomposable algebras of degree $p^{n}$ and exponent $p$ exist ([Kar95]). If the essential dimension were $\leqslant n$, then the algebra would descend to a field of $p$-rank $\leqslant n$, and therefore of symbol length at most $n$. That would mean the algebra decomposes as a tensor product of $\leqslant n$ cyclic algebras of degree $p$. Hence, indecomposable algebras of degree $p^{n}$ and exponent $p$ are of essential dimension at least $n+1$.
@@
However, since the essential dimension of the generic totally decomposable algebras is $n+1$, one would expect that indecomposable algebras are of greater essential dimension.
Question 10.1.
@@
What is the essential dimension of the indecomposable algebras of degree $p^{n}$ and exponent $p$ constructed in [Kar95]?
@@
This lower bound of $n+1$ for $\operatorname{ed}(\operatorname{Alg}_{p^{n},p})$ is the current state of the art. For upper bounds, in [GG16, Theorem 1.3], an upper bound of $2n^{2}-3n-6$ was provided for $\operatorname{ed}(\operatorname{Alg}_{2n,2})$ when $n$ is even, $n>4$ and $\operatorname{char}(k)=2$. Replacing $n$ with $2^{n-1}$ yields $\operatorname{ed}(\operatorname{Alg}_{2^{n},2})\leqslant 2^{2n-1}-3\cdot 2^{n-1}-6$ for $n\geqslant 4$. For results on $\operatorname{Alg}_{4,2}$ and $\operatorname{Alg}_{8,2}$, see Sections 4 and 5. In [Bae12], it was proven that $\operatorname{ed}_{2}(\operatorname{Alg}_{2^{n},2})\leqslant 2^{2n-2}$.
11. Linear Algebraic Groups
Let $m,n$ be positive integers with m|nm\,|\,n. The study of the essential dimension of $\operatorname{Alg}_{n,m}$ is tightly connected to the essential dimension of certain linear algebraic groups. Given that $H^{1}(F,\operatorname{PGL}_{n}(F))$ is bijective to the set of isomorphism classes of central simple $F$-algebras of degree $n$, the essential dimension of $H^{1}(-,\operatorname{PGL}_{n})$ coincides with the essential dimensions of $\operatorname{Alg}_{n,n}$. Now, because of the exact sequence
$1\rightarrow\operatorname{G}_{\text{m}}\rightarrow\operatorname{GL}_{n}/\mu_{m}\rightarrow\operatorname{PGL}_{n}\rightarrow 1,$
$H^{1}(F,\operatorname{GL}_{n}/\mu_{m})$ is bijective to the kernel of the connecting map
$\delta:H^{1}(F,\operatorname{PGL}_{n})\rightarrow H^{2}(F,\operatorname{G}_{\text{m}})=Br(F)$
where each algebra is mapped to the $m$th power of its Brauer class. Thus, $H^{1}(F,\operatorname{GL}_{n}/\mu_{m})$) is bijective to the set of isomorphism classes of central simple $F$-algebras of degree $n$ and exponent dividing $m$ (see [BM09, Example 1.1] for details).
We consequently obtain
$\operatorname{ed}(\operatorname{Alg}_{n,n})=\operatorname{ed}(\operatorname{PGL}_{n}),\ \operatorname{ed}_{p}(\operatorname{Alg}_{n,n})=\operatorname{ed}_{p}(\operatorname{PGL}_{n}),$
$\operatorname{ed}(\operatorname{Alg}_{n,m})=\operatorname{ed}(\operatorname{GL}_{n}/\mu_{m}),\ \operatorname{ed}_{p}(\operatorname{Alg}_{n,m})=\operatorname{ed}_{p}(\operatorname{GL}_{n}/\mu_{m}).$
In [Bae12, Lemma 6.1] and [CM13, Theorem 1.1], it was also proven that
$\operatorname{ed}(\operatorname{Alg}_{n,m})\leqslant\operatorname{ed}(\operatorname{SL}_{n}/\mu_{m})\leqslant\operatorname{ed}(\operatorname{Alg}_{n,m})+1,$
$\operatorname{ed}_{p}(\operatorname{Alg}_{n,m})\leqslant\operatorname{ed}_{p}(\operatorname{SL}_{n}/\mu_{m})\leqslant\operatorname{ed}_{p}(\operatorname{Alg}_{n,m})+1,$
under the assumption that $\operatorname{char}(k)\nmid n$. When $\operatorname{char}(k)|n$, the inequality $\operatorname{ed}(\operatorname{SL}_{n}/\mu_{m})\leqslant\operatorname{ed}(\operatorname{Alg}_{n,m})+1$ still holds true ([Bae12, Remark 1 following Lemma 6.1]).
Question 11.1.
Does $\operatorname{ed}(\operatorname{Alg}_{n,m})\leqslant\operatorname{ed}(\operatorname{SL}_{n}/\mu_{m})$ hold true when $\operatorname{char}(k)=p$?
Via this connection to linear algebraic groups, the following upper bounds are obtained from [GG16]
$\operatorname{ed}(\operatorname{Alg}_{n,n})\leqslant n^{2}-3n+1,\ \text{when}\ n\geqslant 4,\ \text{and}$
$\operatorname{ed}(\operatorname{Alg}_{2n,2})\leqslant 2n^{2}-3n-6,\ \text{when}\ n>4\ \text{and}\ \operatorname{char}(k)|n.$
Acknowledgements
@@
The authors are indebted to Jean-Pierre Tignol and Uzi Vishne for their helpful remarks on an earlier version of the manuscript.

### v2

@@ 8. Algebras of prime degree
Fix a prime degree $p$ and a field $F$ of $\operatorname{char}(F)=p$. Recall a cyclic algebra of degree $p$ over $F$ has a particularly nice presentation as it is generated by two elements $i$ and $j$ such that $i^{p}-i=\alpha$ for some $\alpha\in F$, $j^{p}=\beta$ for some $\beta\in F^{\times}$ and $jij^{-1}=i+1$. This algebra is denoted by the symbol $\alpha,\beta)_{p,F}$. The extension $F[i]/F$ is either a cyclic field extension of degree $p$ with Galois group $\langle\sigma\rangle$ where $\sigma(i)=i+1$ or $F\times\dots\times F$, an étale extension. In both cases, we can define the norm $N_{\alpha}:F[i]\rightarrow F$ by $x\mapsto x\sigma(x)\cdots\sigma^{p-1}(x)$ where $\sigma(i)=i+1$. The algebra $\alpha,\beta)_{p,F}$ is “split”, i.e., isomorphic to $M_{p}(F)$, when $\beta=N_{\alpha}(y)$ for some $y\in F[i]$. If no such $y$ exists, then $\alpha,\beta)_{p,F}$ is a division algebra. Different symbols can represent the same isomorphism class. In particular $\alpha,\beta)_{p,F}=[\alpha+\lambda^{p}-\lambda+x^{p}\beta^{n},\beta)_{p,F}$ for any choice of $\lambda,x\in F$ and $n\in\{1,\dots,p-1\}$, and $\alpha,\beta)_{p,F}=[\alpha,\beta N_{\alpha}(y))_{p,F}$ for any $y\in F[i]$ (note that we use == also for $\cong$ as is common in the literature).
It is still unknown whether every algebra of degree $p$ over $F$ is cyclic. This is known when $p=2$ or $3$, but open for $p\geqslant 5$. In the case of $p$-special fields, however, this is known to be true.
Theorem 8.1 ([1]).
@@
This is an immediate result of Theorem [1, Chapter IV, Theorem 32]. However, we include an alternative proof here using the Nullstellensatz for $p$-special fields.
Consider $A$ of degree $p$ over $F$. A necessary and sufficient condition for $A$ to be cyclic is the inclusion of a non-central element $y$ satisfying $y^{p}\in F$ ([1, Lemma 10]). Recall that there are the characteristic coefficients $f_{1},\dots,f_{p-1},f_{p}$ from $A$ to $F$, and that each $x\in A$ satisfies $x^{p}-f_{1}(x)x^{p-1}-\dots-f_{p-1}(x)x-f_{p}(x)=0$. It is therefore enough to find a nontrivial solution to the system $f_{1}(x)=f_{2}(x)=\dots=f_{p-1}(x)=0$ over a subspace of $A$ consisting of noncentral elements, and then the solution $y$ will be a noncentral element satisfying $y^{p}\in F$. Since the center of $A$ is a subspace of dimension 1, there exists a subspace $V$ (satisfying $A=V\oplus F$) consisting of noncentral elements (apart from zero) of dimension $p^{2}-1$. By the Nullstellensatz for $p$-special fields ([41, Theorem 1.13]), there exists a nontrivial solution to the system $f_{1}(x)=f_{2}(x)=\dots=f_{p-1}(x)=0$ over $V$, and that settles the issue. ∎
The following result is analogous to [44, Lemma 8.5 (7)] which was stated in the good characteristic case:
Corollary 8.2.
The essential $p$-dimension of $\operatorname{Alg}_{p,p}$ is $2$ when the base field $k$ is algebraically closed and of characteristic $p$.
@@
Let $F$ be a field containing $k$. Every cyclic $F$-algebra, $\alpha,\beta)$ is defined over $k(\alpha,\beta)$, a field of transcendence degree at most 2 over the base field, $k$. By the previous theorem, every algebra of degree $p$ over $F$ becomes cyclic under some prime-to-$p$ extension $L/F$, and therefore, the essential $p$-dimension of $\operatorname{Alg}_{p,p}(F)$ is at most 2. In order to show that it is exactly 2, it is enough to point out one cyclic division algebra of degree $p$, because the transcendence degree bounds the cohomological $p$-dimension, and for the existence of division cyclic algebras, it must be at least 2. By [17, Remark] (using valuation theory, see [48] for background), the generic cyclic algebra $\alpha,\beta)_{p,F}$ is division for $F=k(\alpha,\beta)$. Since it remains division under prime-to-$p$ extension, the statement readily follows.
@@ 9. Algebras of Prime Exponent
By [24, Theorem 9.1.4], which dates back to Teichmüller, ${{}_{p}{\operatorname{Br}}}(F)$ is generated by cyclic algebras of degree $p$ for fields $F$ of $\operatorname{char}(F)=p>0$. Tensor products of cyclic algebras of degree $p$ therefore have a special place in the theory of central simple algebras.
@@
As mentioned in Section 7, there is a description of ${{}_{p}{\operatorname{Br}}}(F)$ in terms of differential forms. Recall that $\Omega_{F}$ is the ring of differential forms over $F$, i.e., sums of terms of the form adbadb, where $a,b\in F$, and $d(bc)=cdb+bdc$. The group ${{}_{p}{\operatorname{Br}}}(F)$ is isomorphic to the cokernel of $\wp$, where $\wp:\Omega_{F}\rightarrow\Omega_{F}/dF$ is given by $a\frac{db}{b}\mapsto(a^{p}-a)\frac{db}{b}$. The isomorphism takes $a\frac{db}{b}$ to $a,b)_{p,F}$ and, as the argument in Section 7 with $n=1$ shows, the $p$-rank of $F$ bounds the symbol length of ${{}_{p}{\operatorname{Br}}}(F)$, i.e., the maximal number of cyclic algebras of degree $p$ required in order to express any given class in ${{}_{p}{\operatorname{Br}}}(F)$ is bounded from above by the $p$-rank of $F$.
@@
Consider the functor $\operatorname{Dec}_{p,n}$ mapping each field $F$ containing $k$ to the set of algebras of degree $p^{n}$ that decompose as tensor products of $n$ cyclic algebras of degree $p$ over $F$. The fact that $\operatorname{ed}((\mathbb{Z}/p\mathbb{Z})^{\times n})$ is 1 over fields of characteristic $p$ ([31]) implies that $\operatorname{ed}(\operatorname{Dec}_{p,n})\leqslant n+1$. Here is a direct proof of this fact without using the result on $\operatorname{ed}((\mathbb{Z}/p\mathbb{Z})^{\times n})$:
Proposition 9.1.
The essential dimension of $\operatorname{Dec}_{p,n}$ is bounded from above by $n+1$.
Proof.
In [21] it was shown that every tensor product of $n$ cyclic algebras of degree $p$ can be written as
$[\alpha_{1},\alpha_{2})_{p,F}\otimes[\alpha_{2},\alpha_{3})_{p,F}\otimes\dots\otimes[\alpha_{n},\alpha_{n+1})_{p,F}.$
@@
This observation is based on the rule $\alpha,\beta)_{p,F}=[\alpha(\beta+x^{p})\beta^{-1},\beta+x^{p})_{p,F}$, which means that each $\alpha,\beta)_{p,F}\otimes[\gamma,\delta)_{p,F}$ can be modified into $\alpha(\beta+x^{p})\beta^{-1},\beta+x^{p})_{p,F}\otimes[\gamma+x^{p}-x,\delta)_{p,F}$ where $x=\gamma-\beta$ (i.e., $\gamma+x^{p}-x=\beta+x^{p}$), and the statement follows by induction. ∎
In [36], the generic tensor product of $n$ cyclic algebras of degree $p$ was used to bound the essential dimensions of both $\operatorname{Dec}_{p,n}$ and $\operatorname{Alg}_{p^{n},p}$ from below by $n+1$. For $p=2$ and $n=3$, this improved the lower bound produced in [5] for $\operatorname{Alg}_{8,2}$ from 3 to 4. We recall here briefly the argument made in [20] for $n+1$ being a lower bound for $\operatorname{Alg}_{p^{n},p}$:
Proposition 9.2.
The essential $p$-dimension of $\operatorname{Alg}_{p^{n},p}$ is at least $n+1$.
Proof.
@@
When $n=1$ the essential $p$-dimension is 2 by Corollary 8.2. Suppose $n\geqslant 2$. The case of $p=n=2$ is also easy, see Proposition 3.1. When either $n\geqslant 3$ or $p\geqslant 3$, indecomposable algebras of degree $p^{n}$ and exponent $p$ exist and remain indecomposable after any prime to $p$ extension ([28]). If the $p$-essential dimension were $\leqslant n$, then a prime to $p$ extension of the algebra would descend to a field of $p$-rank $\leqslant n$, and therefore have symbol length at most $n$. That would mean the algebra decomposes as a tensor product of $\leqslant n$ cyclic algebras of degree $p$. Hence, indecomposable algebras of degree $p^{n}$ and exponent $p$ are of $p$-essential dimension at least $n+1$. ∎
@@
Since the essential dimension of the generic totally decomposable algebras is $n+1$, one might expect that indecomposable algebras to be of greater essential dimension.
Question 9.3.
@@
What is the essential dimension of the indecomposable algebras of degree $p^{n}$ and exponent $p$ constructed in [28]?
@@
This lower bound of $n+1$ for $\operatorname{ed}(\operatorname{Alg}_{p^{n},p})$ is the current state of the art. For upper bounds, in [23, Theorem 1.3], an upper bound of $2n^{2}-3n-6$ was provided for $\operatorname{ed}(\operatorname{Alg}_{2n,2})$ when $n$ is even, $n>4$ and $\operatorname{char}(k)=2$. Replacing $n$ with $2^{n-1}$ yields $\operatorname{ed}(\operatorname{Alg}_{2^{n},2})\leqslant 2^{2n-1}-3\cdot 2^{n-1}-6$ for $n\geqslant 4$. For results on $\operatorname{Alg}_{4,2}$ and $\operatorname{Alg}_{8,2}$, see Sections 3 and 4. In [6], it was proven that $\operatorname{ed}_{2}(\operatorname{Alg}_{2^{n},2})\leqslant 2^{2n-2}$.
Acknowledgments
@@
The authors are greatly indebted to Jean-Pierre Tignol, Uzi Vishne, and the anonymous referees for their helpful remarks on earlier versions of the manuscript.